OpenStudy (anonymous):

What are the solutions to Ix-10I -4=2x?

OpenStudy (anonymous):

@Vocaloid

OpenStudy (hhopke):

x=6

Vocaloid (vocaloid):

alright, this one is a bit trickier first, let's add 4 to each side to get the absolute value by itself |x-10| = 2x + 4 now we write two equations: x - 10 = 2x + 4 x - 10 = -(2x+4) solve both for x. you will get 2 different values

OpenStudy (anonymous):

x=2 and x=-14?

Vocaloid (vocaloid):

good, now plug both back into the original equation. only one ends up working in the end

OpenStudy (anonymous):

wait absolute value means that the - is actually positive unless written like this -|10|

OpenStudy (anonymous):

x=2?

Vocaloid (vocaloid):

good, x = 2 is our answer. x = -14 is an extraneous solution

OpenStudy (anonymous):

wouldn't it be 3 ?????????????????????????

OpenStudy (anonymous):

everything that is enclosed in an absolute bars | | equals to zero an unknown number with variable x may either be positive or negative |-x| = x and |x| = x this is why we have two sets of answers for the most part

Vocaloid (vocaloid):

or, to do it algebraically: x - 10 = -(2x+4) x - 10 = -2x - 4 3x = 6 x = 2

OpenStudy (anonymous):

in this case wouldnt it be x+10 and |x+(-10)| be the same or no and why?

Vocaloid (vocaloid):

x = 3 doesn't work, try plugging it back into the original equation

Vocaloid (vocaloid):

no, x+10 and |x+(-10)| are not the same

OpenStudy (anonymous):

o yh your right I plugged 2 in sorry

OpenStudy (anonymous):

I got it ;)

OpenStudy (anonymous):

AND X Does Not Equal 6 -_-